a manufacturer of fluorescent bulbs claims that mean life of fluorescent bulb is 400 hours. a sample of 35 fluorescent bulbs was taken and it showed a mean of 399 hours with a standard deviation of 1.1 hours, is the mean life different from 400 hours? use the 0.05 significance level.
1
Expert's answer
2021-07-19T05:42:57-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=400"

"H_1:\mu\not=400"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05, df=n-1"

"=35-1=34" ​​degrees of freedom, and the critical value for a two-tailed test i"t_c=2.032244."

The rejection region for this two-tailed test is "R=\{t:|t|>2.032244\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{399-400}{1.1/\sqrt{35}}=-5.378254"

Since it is observed that "|t|=5.378254>2.032244=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for two-tailed "\alpha=0.05, df=34,"

"t=-5.378254," ​is"p=0.000006," ​and since"p=0.000006<0.05=\alpha," it is concluded that the null hypothesis is rejecte

Therefore, there is enough evidence to claim that the population mean "\mu" is different than "400," at the "\alpha=0.05" significance level.